OpenStudy (anonymous):

Find the area for each circle with a diameter of 11ft. A. 90 ft2 B. 95 ft2 C. 97 ft2 D. 99 ft2

jimthompson5910 (jim_thompson5910):

If the diameter is 11 ft, then the radius is half that at r = 11/2 = 5.5 feet Now plug this into the formula A = pi*r^2 A = pi*r^2 A = pi*(5.5)^2 A = (3.14)*(5.5)^2 I'll let you finish

OpenStudy (smarty):

That doesn't seem right... how many circle are there?

OpenStudy (smarty):

the formula and set up is correct... just multiply your answer by however many circles there are to get the final answer.

OpenStudy (anonymous):

There iss no picture thoee o; . . thats thee whole problemm .

OpenStudy (smarty):

Hm... cuz in your question it says find the area for EACH circle.... :/ With the given diameter and set up, your answer should be 30.25... I would go with A. Because you can assume that there are 3 circles... with is appx. 90.75 thus making the other options invalid

jimthompson5910 (jim_thompson5910):

not sure how you're getting 30.25 since (3.14)*(5.5)^2 = 94.985

OpenStudy (anonymous):

Isnt 95 feet ?

OpenStudy (smarty):

My bad, you should go with B, my error i forgot to multiple pi.

jimthompson5910 (jim_thompson5910):

yes so it's closest to 95 square feet

OpenStudy (anonymous):

if you round too the wholee number ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thank youuu (:

jimthompson5910 (jim_thompson5910):

yw